OpenStudy (anonymous):

HELP FAST Which equation is a parabola with a focus of (2,2) and a directrix of Y=4.

OpenStudy (aum):

If (x,y) is any point on the parabola, then it will be equidistant from the focus and the directrix. (x-2)^2 + (y-2)^2 = (y-4)^2 (x-2)^2 = (y-4)^2 - (y-2)^2 = (y-4+y-2) * (y-4-y+2) = (2y-6)(-2) = -4(y - 3) \(\Large y = -\frac 14(x-2)^2 + 3 \)

OpenStudy (anonymous):

TY!!!! @aum

OpenStudy (aum):

You are welcome.